In algebra when adding a negative number, does it mean the same thing as subtracting a positive number?
Found 2 solutions by ewatrrr, rfer:
Answer by ewatrrr(24785)


Hi,
adding a negative number, for ex:
7 + (-4) = 7 - 4 = 3
Hope this answers your question

Answer by rfer(16322)

Look at the number line, start from the first number then move more negative by the value of the negative number.
Your statement works if it is a positive number and a negative number, but if they are both negative numbers, they add making the answer more negative.
Does that help?
Bob